# HealthQ: Unveiling Questioning Capabilities of LLM Chains in Healthcare Conversations

### v3

@@ 3.4 Evaluation Framework
Table 3: Comparison of Summary-Based Metrics across LLM Chains
(a) Information Extraction Quality - NER2NER
LLM Healthcare Chain Medication (%) Symptoms (%) Total (%)
Hardcoded 0.0022 0.0384 0.0317
RAG_default 0.0167 0.0942 0.0727
RAG 0.0126 0.0836 0.0777
RAG_reflection 0.0426 0.1434 0.1350
RAG_reflection_CoT 0.0222 0.1054 0.0960
RAG_reflection_CoT_SC 0.0280 0.0825 0.0720
ReAct 0.0000 0.0072 0.0066
(b) Summarization Performance - ROUGE
LLM Healthcare Chain rouge-1 F1 rouge-2 Recall rouge-2 Precision rouge-2 F1 rouge-L Recall rouge-L Precision rouge-L F1
Hardcoded 0.0382 0.0000 0.0000 0.0000 0.0280 0.0647 0.0369
RAG_default 0.1426 0.0241 0.0156 0.0164 0.1853 0.1119 0.1262
RAG 0.0991 0.0089 0.0170 0.0106 0.0716 0.1290 0.0861
RAG_reflection 0.1931 0.0652 0.0459 0.0480 0.2077 0.1649 0.1675
RAG_reflection_CoT 0.1447 0.0272 0.0245 0.0216 0.1337 0.1430 0.1230
RAG_reflection_CoT_SC 0.1275 0.0170 0.0130 0.0128 0.1243 0.1125 0.1064
ReAct 0.0840 0.0065 0.0049 0.0046 0.0856 0.0840 0.0756
3.5 Validation
To ensure the validity of our evaluation, we devised a framework distinct from traditional interrogation or cross-validation methods. Our approach does not directly rely on ground truth from a test set to evaluate the system. Instead, it employs separate LLM workflows for the roles of judge and doctor. The judge LLM, which has access to the test ground truth data, is responsible for evaluating the questions generated by the LLM healthcare chains. In contrast, the LLM healthcare chains, acting as doctor LLM, only have access to the virtual patient initial statement $Stmt$. This separation ensures that the question generation process remains unbiased and prevents data leakage.
To maintain the integrity of the evaluation, we strictly prevent data leakage. The LLM healthcare chains never directly access the test patient known knowledge $K$. Thus, the phrasing of the questions is based solely on a subset of $K$ and the search results from the chains’ own knowledge base. This design ensures that the evaluation of the generated questions is impartial and based on the information provided in the virtual patient statement.
HealthQ addresses two critical research questions (RQs) to validate its effectiveness:
  • 1.
    RQ1: How well does the induced answer to the question summarize all the relevant information? This question assesses the completeness and quality of the answers generated by the LLM virtual patient in response to their own questions.
  • 2.
    RQ2: How are the results of LLM judge interrogation for the proposed questions correlated with the quality of the induced answers? This question explores the relationship between the quality of the questions and the quality of the answers they elicit, using well-established summarization metrics.
Table 4: LLM Judge Interrogation Results
LLM Healthcare Chain Specificity Usefulness Relevance Coverage Fluency
Hardcoded 3.56 4.52 4.66 2.30 5.53
RAG_default 5.30 5.95 6.50 3.58 8.06
RAG 7.78 7.87 8.75 4.55 9.16
RAG_reflection 7.78 7.94 8.96 5.08 8.92
RAG_reflection_CoT 7.84 8.08 8.98 5.02 9.02
RAG_reflection_CoT_SC 7.93 8.44 9.29 5.34 8.88
ReAct 3.55 4.65 3.46 2.26 7.59
For RQ1, we evaluate the effectiveness of the questions using a virtual patient and established summarization metrics, measuring how well the induced answers encapsulate patient information. For RQ2, we analyze the correlation and mutual information between the quality of the questions and the comprehensiveness of the answers. A strong correlation would suggest that well-formulated questions lead to more informative answers, thereby enhancing patient information gathering and improving healthcare outcomes.
Mutual information $I(X_{i},X_{j})$ quantifies shared information between variables without the need of predictive modeling. By using its standardized form $\frac{I(X_{i},X_{j})}{\min(H(X_{i}),H(X_{j}))}$, where $H(X_{i})$ is the entropy of $X_{i}$ and $H(X_{j})$ is the entropy of $X_{j}$, we identify dependencies or shared information within the interrogation and summarization groups. This ratio indicates the proportion of shared information relative to the smaller entropy value, providing a normalized measure of pairwise dependence. Although discretizing data into bins may introduce numerical errors, the coarsened scores can still effectively achieve the evaluation goal, helping us understand the inherent relationships in our evaluation.
Figure 2: LLM judge interrogation metrics: kernel density estimation and radar plots
Our mutual information validation comprises two steps:
  1. 1.
    Individual Data Point Analysis: This analysis examines the relationship between the interrogation score for each individual question generated by the LLM healthcare chains and the informativeness of the corresponding response from the virtual patient. Specifically, it assesses whether a high score in interrogation metrics (see Table 2) for a single question leads to a more comprehensive and informative answer for that specific interaction.
  2. 2.
    Aggregate Analysis: This analysis evaluates the overall performance of each LLM healthcare chain by averaging the interrogation scores across all interactions. It then investigates whether the average interrogation score for a given LLM healthcare chain correlates with the overall quality and comprehensiveness of the answers induced by that chain across the entire dataset. This helps in understanding whether LLM chains that consistently generate high-quality questions also tend to elicit more informative responses in general.
The results of these analyses, presented through correlation and mutual information matrices in Section 4, provide insights into how well-formulated questions can enhance patient interactions and improve healthcare outcomes.
4 Experiments
4.1 Experiment Setup
To ensure the feasibility of our approach for local implementations in clinics and hospitals, all components of the LLM health chains are based on open-source models. The primary LLM used is the Mixtral [^8] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 8X7b, served on Groq [^9] https://wow.groq.com/. On branches of Chain of Thought Self-Consistency (CoT-SC) processes, a temperature of 0.3 is used, while a temperature of 0 is applied for all other models. The NER component utilizes an open-source biomedical NER model [35] to extract medical entities to make set comparisons. We employ FAISS [36] with its default settings as the VDB index package, and the MiniLM [^10] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 model for text embeddings. Data preprocessing—including the extraction of patient known knowledge ($K$), virtual patient simulation, and question evaluation by the LLM judge—is conducted using Claude-3-Opus [^11] https://claude.ai/. The experiments are conducted on a dataset comprising 128 test cases (64 from each of the two datasets). We benchmark seven different LLM health chains, generating a total of 796 patient status-doctor statement pairs. For comparative analysis, we group the data by the LLM chain type and summarize the mean scores in Table 4.
(a) Individual interactions
(b) Aggregated within specific LLM healthcare chains
Figure 3: Mutual information analysis - correlation between question quality and answer quality, across all metrics
This setup rigorously evaluates the questioning capabilities of various LLM healthcare chains, ensuring comprehensive and reliable results.
4.2 Results and Analysis
@@
The summary-based metrics presented in Table 3(b) are particularly relevant to RQ1, which assesses how well the induced answers summarize all relevant information. The results show that state-of-the-art chains like RAG_reflection and RAG_Reflection_CoT_SC consistently outperform more primitive methods, such as the hardcoded chain, across nearly all evaluated metrics. This is especially evident in the ROUGE and NER2NER metrics, where these advanced chains demonstrate superior performance. For instance, the RAG_reflection_CoT_SC chain achieves the highest overall scores in several categories, reflecting its ability to generate contextually relevant and informative questions that lead to more comprehensive and accurate patient responses. The lower performance of the ReAct chain could be attributed to its design in LangChain, emphasizing answering rather than questioning, along with the limited number of tools provided to it. These findings suggest that well-formulated questions by the more advanced chains lead to better information extraction, thereby validating the effectiveness of our evaluation framework in summarizing all relevant information, as addressed in RQ1.
For RQ2, which examines the correlation between the quality of questions generated by LLM chains and the resulting answers, the LLM judge interrogation results in Table 4 show that most LLM healthcare chains substantially outperform the hardcoded workflow across all interrogation metrics. The basic RAG chain shows notable improvements in specificity, usefulness, relevance, and fluency compared to the hardcoded method, highlighting its more flexible and adaptive approach to generating questions. Chains that incorporate reflection and iterative reasoning techniques, such as those using reflection and CoT, exhibit further enhancements, particularly in relevance and coverage. The reflection mechanism enables the chains to reassess and refine the context of the questions, leading to more targeted and contextually appropriate queries. The addition of self-consistency checks further boosts performance, with the RAG_CoT_SC chain showing the highest scores in relevance and fluency. This improvement is attributed to the iterative reasoning and self-consistency processes, which help in formulating more coherent and relevant questions. The reaction-based chain, while showing some improvements over the hardcoded method, still lags behind the more advanced RAG variants, particularly in specificity and coverage. This analysis demonstrates that chains performing well in interrogation metrics tend to produce better quality answers, thus validating the connection explored in RQ2.
Figure 2 presents the kernel density estimation (KDE) and radar plots comparing the distributions of interrogation metrics across LLM healthcare chains, providing further insights into RQ2. The KDE plots reveal that advanced chains, such as RAG and RAG with reflection, exhibit broader and higher-density distributions in key metrics like specificity, relevance, and usefulness. This reflects their adaptive questioning capabilities and iterative reasoning processes. Chains incorporating reflection and CoT techniques show distinct improvements, particularly in relevance and specificity, due to their ability to refine questions based on previous outputs. The radar plot consolidates these findings, with self-consistency-enhanced chains (RAG_reflection_CoT_SC) displaying more balanced performance across all metrics. In contrast, the hardcoded chain has narrow distributions, indicating rigid question generation, while the ReAct chain struggles with fluency and coverage, possibly due to limitations in its toolset. These observations support the idea that well-designed chains, especially those employing reflection and CoT methods, are more effective in generating questions that lead to informative patient responses, aligning with the goals outlined in RQ2.
Figures 3(a) and 3(b) offer deeper insights into the relationship between question quality and answer quality using normalized mutual information (MI) matrices. In Figure 3(a), MI is computed between two metrics, $J_{m}$ and $J_{n}$, for each individual interaction:
$MI\left(J_{m}(data_{p}),J_{n}(data_{p})\right)$
where $p$ indexes individual interactions or patient cases. This figure shows strong dependence between metrics such as ”Usefulness” and ”Relevance” ($MI=0.50$), reflecting their shared focus on capturing relevant aspects of the patient’s condition. In contrast, pairs with less dependency such as those between ”Fluency” and ”Coverage” (MI = 0.38) indicate weaker relevance. In this case, fluency does not necessarily imply the completeness of aspects covered in the question. In Figure 3(b), MI is calculated based on aggregated scores across all interactions within each LLM chain type:
$MI\left(J_{m}\left(\sum_{i\in Chain[l]}data_{i}\right),J_{n}\left(\sum_{i\in Chain[l]}data_{i}\right)\right)$
Here, the summation $\sum_{i\in Chain[l]}data_{i}$ represents all interactions $i$ for a specific chain type $Chain[l]$. This aggregated analysis reveals that correlations within each chain type remain consistent, with strong relationships among metrics like ”Usefulness” and ”Relevance.” Chains that score well in these areas also tend to perform well in summarization metrics like ROUGE and NER2NER, affirming the robustness of the evaluation framework.
Overall, the results indicate that higher interrogation scores correlate with better summarization performance, confirming our hypothesis that good questions lead to better information extraction. This correlation underscores the importance of effective questioning in improving the overall quality of patient interactions and diagnostic outcomes, thereby demonstrating the value of our evaluation framework.
5 Discussion
5.1 Sources of Bias and Correction

### v4

@@ III-D Evaluation Framework
III-E Validation
Our validation framework ensures the robustness and reliability of the evaluation by focusing on the dependencies between interrogation and summarization metrics. The analysis is conducted at an aggregate level, where the scores for interrogation and summarization metrics are averaged across all interactions for each LLM healthcare chain. This approach captures systematic patterns and relationships across chain implementations, minimizing the influence of isolated cases or outliers.
To quantify these dependencies, we employ normalized mutual information (NMI), which measures the shared dependency between two metrics. Mutual information $I(X,Y)$ evaluates how much knowing one variable reduces uncertainty about the other:
$I(X,Y)=\sum_{x\in X}\sum_{y\in Y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)},$
where $p(x,y)$ represents the joint probability distribution of $X$ and $Y$, while $p(x)$ and $p(y)$ are their respective marginal probabilities. Mutual information captures nonlinear dependencies and provides a general measure of association between the metrics.
Since the magnitude of raw mutual information depends on the scales of $X$ and $Y$, we normalize it to ensure comparability across different metrics. The normalized mutual information (NMI) is computed as:
$NMI(X,Y)=\frac{I(X,Y)}{\min(H(X),H(Y))},$
where $H(X)$ and $H(Y)$ denote the entropies of $X$ and $Y$, respectively:
$H(X)=-\sum_{x\in X}p(x)\log p(x).$
NMI values range from 0 to 1, where 0 indicates no dependency, and 1 signifies a perfect association between $X$ and $Y$. For example, a high NMI between interrogation metrics such as Relevance and summarization metrics like ROUGE-L Recall indicates that questions rated as highly relevant are more likely to elicit comprehensive answers aligning with the ground truth. By focusing on NMI, we capture both linear and nonlinear relationships between the metrics, offering a holistic view of their dependencies.
The aggregate analysis calculates average scores for each interrogation and summarization metric across all interactions within a given LLM healthcare chain. This averaging process ensures that the dependencies identified reflect the overall performance of the chain rather than being skewed by individual instances. Figure 3 illustrates the NMI matrix for the aggregated metrics, providing a comprehensive view of the relationships between interrogation and summarization metrics across the LLM healthcare chains.
The validation process is designed to systematically address the core research questions:
  • RQ1: How well does the induced answer to the question summarize all the relevant information? By analyzing the dependencies between summarization metrics (e.g., ROUGE, NER2NER) and interrogation metrics, we assess whether high-quality questions generate responses that effectively encapsulate the ground truth.
  • RQ2: How correlated are the LLM judge interrogation metrics with the quality of the induced answers? By leveraging NMI to quantify dependencies, we determine whether chains generating better-rated questions (e.g., higher Relevance or Usefulness) also elicit more comprehensive and informative answers.
By conducting the analysis at the aggregate level, we ensure that the framework captures general trends and patterns across different chains, reducing the noise introduced by outlier interactions. Furthermore, the use of normalized mutual information ensures that the evaluation is consistent and comparable across diverse metrics. The structure of the validation framework, which separates question generation from evaluation and prevents direct access to test ground truth, ensures fairness, avoids data leakage, and strengthens the credibility of the evaluation process. To facilitate better visualization and comparison across metrics with varying scales, scores were normalized to a common range through a linear rescaling process. This ensures consistent representation of the performance metrics while preserving the relative differences between them.
IV Experiments
IV-A Experiment Setup
To ensure the feasibility of our approach for local implementations in clinics and hospitals, all components of the LLM health chains are based on open-source models where possible. The primary LLM used for the health chain processes is Mixtral [^5] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1, served on Groq [^6] https://wow.groq.com/. For evaluating the generated questions, GPT-3.5-turbo and GPT-4 (via OpenAI API [^7] https://platform.openai.com/docs/models/gpt-4), along with Claude-3-Opus [^8] https://claude.ai/, are used as LLM judges. These models ensure robustness in evaluation and enable comparative analysis across different judging capabilities.
For chains using Chain of Thought Self-Consistency (CoT-SC) processes, a temperature of 0.3 is used to introduce slight variability, while a temperature of 0 is applied for all other models to ensure deterministic outputs. The NER component leverages an open-source biomedical NER model [41] to extract medical entities, which are then compared using set-based similarity metrics.
For retrieval tasks, we employ FAISS [42] with its default settings as the vector database (VDB) indexing package. Text embeddings are generated using the MiniLM [^9] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 model. Data preprocessing—including the extraction of patient known knowledge ($K$), virtual patient simulation, and question evaluation by the LLM judge—is conducted using Claude-3-Opus, GPT-3.5-turbo, and GPT-4 to ensure consistency and coverage in the evaluation pipeline.
The experiments are conducted on a dataset comprising 128 test cases (64 from each of the two datasets). We benchmark seven different LLM health chains, generating a total of 796 patient status-doctor statement pairs. For comparative analysis, the data is grouped by LLM chain type, and evaluation results are summarized for each model used as a judge. This includes mean scores from interrogation metrics such as specificity, usefulness, relevance, coverage, and fluency (see Table III). The results are further validated by analyzing correlations between the LLM judges and traditional metrics like ROUGE and NER2NER.
IV-B Evaluation Metrics Implementation
To robustly evaluate the ability of LLM healthcare chains to gather useful patient information through effective questioning, we employ three complementary approaches: LLM judge interrogation, NER-based comparison, and ROUGE-based summarization evaluation.
The LLM judge interrogation assesses question quality using five key metrics: Relevance, Usefulness, Specificity, Coverage, and Fluency. These metrics evaluate how well questions target essential patient information, their potential to uncover diagnostic insights, alignment with patient context, comprehensiveness, and linguistic clarity. Scores range from 0 to 1, with higher scores indicating better performance. Evaluations are conducted using advanced models such as GPT-3.5, GPT-4, and Claude-3-Opus to ensure robustness and diversity.
The NER-based comparison leverages a biomedical NER model [41] trained to identify key medical entities such as symptoms (e.g., physical signs or complaints) and medications (e.g., treatments or prescriptions). It evaluates how effectively the generated questions elicit responses containing relevant medical information by comparing the entities extracted from answers against the ground truth patient data.
The ROUGE-based summarization evaluation measures how closely the information elicited by the LLM-generated questions aligns with the comprehensive ground truth patient information. ROUGE-1 examines word-level overlap for basic coverage, ROUGE-2 assesses two-word phrase overlap to capture more complex patterns, and ROUGE-L evaluates the longest matching subsequences to gauge overall flow and completeness.
The gold standard for all evaluations is derived from the original patient notes in the test set, encompassing a complete range of symptoms, medical history, and contextual details. This comprehensive ground truth ensures that our framework rigorously measures both the breadth and depth of patient information elicited by the questions. Together, the LLM judge, NER, and ROUGE metrics quantify the relationship between question quality and the informativeness of elicited answers, providing a holistic assessment of the LLM healthcare chains’ effectiveness.
IV-C Results and Analysis
TABLE III: Performance of LLM Healthcare Chains Evaluated by LLM Judge Metrics
LLM Chain GPT-4o GPT-3.5 Claude
Specificity Usefulness Relevance Coverage Fluency Specificity Usefulness Relevance Coverage Fluency Specificity Usefulness Relevance Coverage Fluency
RAG 7.38 8.27 8.19 4.54 9.36 7.18 7.99 8.98 5.18 9.08 7.31 7.96 8.88 4.89 9.06
RAG_default 6.80 7.95 7.89 4.56 9.01 6.30 7.20 8.59 4.94 9.00 5.93 6.07 6.69 5.73 8.24
RAG_reflection 7.98 8.82 8.72 5.39 9.47 7.84 8.53 9.18 5.77 9.23 7.88 8.38 9.28 5.64 9.17
RAG_reflection_cot 7.68 8.54 8.43 5.11 9.33 7.50 8.21 9.05 5.53 9.13 7.57 8.13 9.07 5.35 9.12
RAG_reflection_cot_sc 7.28 8.12 8.03 4.90 9.12 7.02 7.89 8.82 5.24 9.00 7.12 7.86 8.79 5.09 8.98
ReAct 5.52 5.59 5.46 3.63 8.44 4.59 5.75 5.68 4.34 7.92 3.65 4.70 3.83 2.29 7.84
Hardcoded 5.30 5.02 4.63 3.69 6.70 5.26 5.18 4.83 4.38 6.80 4.30 5.31 5.41 3.70 6.57
Figure 2: LLM judge interrogation metrics: radar plots
(a) NER Metrics Evaluation
Model Chain NER2NER_sym NER2NER_med NER2NER_total
GPT-4o RAG 0.37 0.27 0.47
RAG_default 0.20 0.15 0.34
RAG_reflection 0.52 0.42 0.65
RAG_reflection_cot 0.44 0.33 0.54
RAG_reflection_cot_sc 0.29 0.23 0.40
ReAct 0.10 0.09 0.15
Hardcoded 0.08 0.07 0.11
GPT-3.5 RAG 0.32 0.24 0.42
RAG_default 0.18 0.14 0.30
RAG_reflection 0.45 0.35 0.57
RAG_reflection_cot 0.38 0.28 0.49
RAG_reflection_cot_sc 0.26 0.20 0.34
ReAct 0.12 0.10 0.17
Hardcoded 0.09 0.07 0.13
Claude RAG 0.34 0.25 0.45
RAG_default 0.22 0.16 0.36
RAG_reflection 0.50 0.40 0.63
RAG_reflection_cot 0.42 0.33 0.54
RAG_reflection_cot_sc 0.31 0.24 0.41
ReAct 0.13 0.10 0.16
Hardcoded 0.10 0.08 0.14
(b) ROUGE Metrics Evaluation
Model Chain rouge-1_f rouge-2_r rouge-2_p rouge-2_f rouge-l_r rouge-l_p rouge-l_f
GPT-4o RAG 0.66 0.44 0.40 0.47 0.62 0.50 0.54
RAG_default 0.60 0.37 0.34 0.41 0.55 0.44 0.47
RAG_reflection 0.75 0.52 0.48 0.57 0.70 0.58 0.62
RAG_reflection_cot 0.68 0.45 0.42 0.49 0.64 0.53 0.56
RAG_reflection_cot_sc 0.48 0.32 0.30 0.34 0.45 0.36 0.39
ReAct 0.28 0.18 0.16 0.20 0.26 0.21 0.23
Hardcoded 0.17 0.12 0.10 0.14 0.15 0.12 0.13
GPT-3.5 RAG 0.70 0.46 0.42 0.50 0.65 0.54 0.57
RAG_default 0.57 0.38 0.35 0.43 0.53 0.44 0.47
RAG_reflection 0.78 0.55 0.51 0.60 0.72 0.61 0.65
RAG_reflection_cot 0.71 0.48 0.44 0.53 0.67 0.56 0.60
RAG_reflection_cot_sc 0.49 0.33 0.31 0.38 0.46 0.37 0.40
ReAct 0.29 0.19 0.17 0.21 0.27 0.22 0.24
Hardcoded 0.18 0.13 0.11 0.15 0.16 0.13 0.14
Claude RAG 0.64 0.43 0.39 0.46 0.60 0.48 0.52
RAG_default 0.55 0.36 0.33 0.39 0.51 0.42 0.45
RAG_reflection 0.72 0.51 0.47 0.55 0.67 0.55 0.59
RAG_reflection_cot 0.65 0.44 0.40 0.48 0.61 0.50 0.54
RAG_reflection_cot_sc 0.46 0.31 0.29 0.33 0.43 0.34 0.37
ReAct 0.26 0.17 0.15 0.19 0.25 0.20 0.22
Hardcoded 0.15 0.11 0.09 0.13 0.14 0.11 0.12
Yellow: Best chain (Reflection), Blue: Second-best (Reflection_CoT), Green: Third-best (RAG)
TABLE IV: (a) NER Metrics: Evaluates entity recognition to assess the information elicited by questions. (b) ROUGE Metrics: Analyzes response comprehensiveness and precision. Together, these metrics demonstrate how effective questions drive informative answers in LLM healthcare chains.
@@ IV-C Results and Analysis
The results in Table IV strongly validate the effectiveness of our framework in addressing RQ1, which assesses how well the induced answers summarize all relevant information. Advanced chains such as RAG_Reflection and RAG_Reflection_CoT consistently outperform simpler baselines like hardcoded and ReAct across all evaluation metrics, including both NER-based and ROUGE scores. For example, RAG_Reflection achieves the highest NER2NER_total scores (0.65 for GPT-4o, 0.57 for GPT-3.5, and 0.63 for Claude), indicating its ability to extract critical medical entities effectively. Similarly, its superior ROUGE-L F scores (0.62, 0.65, and 0.59, respectively) highlight its capacity to generate responses that comprehensively align with ground truth patient information. Importantly, the consistency of these results across diverse LLM judges—GPT-4o, GPT-3.5, and Claude—demonstrates the robustness and model-agnostic nature of our framework. Despite the intrinsic differences in these LLMs, our evaluation framework consistently identifies similar performance trends, reinforcing its reliability as a tool for assessing questioning capabilities. Moreover, the significant performance gap between advanced chains and simpler baselines underscores the importance of well-crafted questions in eliciting informative responses. For instance, the hardcoded chain’s lower ROUGE-L F scores (0.13 across judges) reveal its inability to generate effective interactions, affirming that high-quality questions play a pivotal role in summarizing patient-relevant information. These findings not only validate the utility of our framework but also emphasize the critical link between questioning capabilities and effective information elicitation, addressing RQ1 comprehensively.
The results in Table III and Fig. 2 provide comprehensive insights into RQ2, which explores the relationship between question quality and answer quality. Advanced LLM healthcare chains such as RAG_reflection and RAG_reflection_CoT demonstrate significantly better performance across interrogation metrics compared to simpler methods like ReAct and hardcoded chains. For instance, RAG_reflection achieves the highest scores in specificity (7.98 for GPT-4o, 7.84 for GPT-3.5, and 7.88 for Claude), reflecting its ability to generate precise and diagnostically focused questions. Similarly, RAG_reflection_CoT shows strong fluency scores (9.33 for GPT-4o, 9.13 for GPT-3.5, and 9.12 for Claude), benefiting from iterative chain-of-thought reasoning that enhances clarity and coherence. Coverage also improves with advanced chains; RAG_reflection achieves a score of 5.39 under GPT-4o, significantly higher than the hardcoded chain’s 3.69. These trends indicate that reflection and iterative reasoning mechanisms enable chains to generate contextually rich and useful questions, leading to higher-quality responses. Conversely, the hardcoded and ReAct chains consistently lag across all metrics, with ReAct scoring as low as 5.46 in relevance and 3.63 in coverage under GPT-4o, underscoring their inability to adapt dynamically to patient contexts. These results confirm that advanced techniques not only improve question quality but also drive more comprehensive patient information elicitation.
Fig. 2 further illustrates the robustness and consistency of the proposed framework across different LLM judge models, including GPT-4o, GPT-3.5, and Claude. The radar plots highlight that advanced chains like RAG_reflection and RAG_reflection_CoT consistently cover larger, more balanced areas, excelling across metrics such as specificity, relevance, and fluency. For example, the fluency of RAG_reflection remains high across all judges (9.47 for GPT-4o, 9.23 for GPT-3.5, and 9.17 for Claude), suggesting that its question-generation process is resilient to the underlying LLM used for evaluation. The ReAct and hardcoded chains, in contrast, show narrow and uneven distributions, reflecting their lack of versatility in generating effective questions. Importantly, the framework’s consistency across judges demonstrates its model-agnostic nature, reinforcing its reliability for assessing questioning capabilities regardless of the LLM base model. Furthermore, the alignment of interrogation metrics with traditional measures, as shown in Table IV, validates the utility of the framework in systematically capturing essential aspects of questioning quality. For instance, RAG_reflection achieves the highest ROUGE-L scores across all judges (0.70 for GPT-4o, 0.72 for GPT-3.5, and 0.67 for Claude), closely correlating with its high relevance and specificity metrics. These findings substantiate the central hypothesis that well-crafted questions lead to better patient responses, demonstrating the framework’s potential for evaluating and optimizing questioning capabilities systematically.
Figure 3 illustrates the relationships between evaluation metrics, showcasing the dependencies between question quality and answer quality across aggregated LLM healthcare chain interactions. Using normalized mutual information (NMI), the analysis quantifies the degree to which metrics such as Relevance, Usefulness, and traditional summarization measures like ROUGE and NER2NER align in evaluating chain performance.
The matrix reveals strong dependencies between Relevance and ROUGE-L Recall ($NMI=0.85$) as well as Usefulness and NER2NER_total ($NMI=0.84$). These findings demonstrate that questions rated highly for their relevance and diagnostic utility are likely to elicit responses that closely align with the ground truth. This underscores the ability of well-targeted questions to uncover critical patient information effectively, reinforcing the connection between LLM-specific interrogation metrics and traditional summarization measures.
In contrast, weaker dependencies are observed between Fluency and NER2NER_med ($NMI=0.66$), as well as Coverage and ROUGE-2 Precision ($NMI=0.53$). While fluency ensures linguistic clarity and coherence, these results indicate that it does not necessarily correlate with the extraction of detailed medical entities or precise information. Similarly, the modest correlation between Coverage and ROUGE-2 Precision suggests that broad or exhaustive questions do not always translate into precise and specific responses, highlighting potential trade-offs between general coverage and specificity.
Specificity, on the other hand, shows a stronger alignment with summarization metrics such as ROUGE-1 Recall ($NMI=0.79$) and NER2NER_total ($NMI=0.71$). This highlights the critical role of specificity in driving targeted and informative interactions. Specific questions are more likely to elicit detailed and relevant answers that span the full spectrum of patient-relevant information, emphasizing their importance in diagnostic contexts.
Figure 3: Aggregated within specific LLM healthcare chains: Mutual information analysis - correlation between question quality and answer quality across all metrics.
Overall, the strong correlations between interrogation metrics like Relevance and Usefulness with summarization measures such as ROUGE-L validate the robustness of the evaluation framework. These results confirm that high-quality, targeted questions consistently lead to more comprehensive and accurate responses. Conversely, the weaker correlations for Fluency emphasize that while clarity is necessary, it does not guarantee informativeness or precision, pointing to the need for a balanced approach in designing LLM healthcare chains. These insights affirm the framework’s ability to capture nuanced dependencies and provide a holistic assessment of question-answering performance.
V Discussion
V-A Sources of Bias and Correction